# Exact computation of Transfer Entropy with Path Weight Sampling

### v1

@@ Abstract
We plot in Fig. 3c and d transfer entropy rates for the two models. Each data point is obtained from steady state trajectories using TE-PWS. In model A, increasing the feedback strength keeps $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ unchanged while monotonically increasing $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$. The former is an empirical result that we expect to formally hold for an OU process, though we have not analytically proven it. The latter arises because $X_{1}$ becomes increasingly correlated with $X_{2}$ and the information is copied accurately by $X_{3}$. In contrast, in model B, both $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$ peak near the switching regime, phenomenologically similar to how a bistable three-node motif with feedback has been shown to behave through approximate theory10. The values of the transfer entropy rates in model B are also amplified manyfold compared to model A due to stronger correlations between $X_{1}$ and $X_{2}$ resulting from the switching.
@@
In Fig. 3e we have plotted the amount of DPI violation in both models quantified as a ratio of transfer entropy rates $\mathcal{T}^{*}=\dot{\mathcal{T}}_{X_{1}\to X_{3}}/\dot{\mathcal{T}}_{X_{1}\to X_{2}}$, as a function of the ratio of the feedback to feedforward strengths, $f^{*}$. Surprisingly, we find that regardless of the nature of the variation of the individual transfer entropy rates with increasing feedback, i.e., monotonic or non-monotonic (Figs. 3c and d), the ratio $\mathcal{T}^{*}$ monotonically increases with increasing feedback (Fig. 3 e). Moreover, the ratio overcomes the DPI bound when the strength of the feedback becomes comparable to that of the feedforward coupling. Our results thus show that when the feedback $X_{2}\to X_{1}$ dominates over the feedforward interaction, $X_{1}\to X_{2}$, the feedforward entropy $\mathcal{T}_{X_{1}\to X_{3}}$ becomes larger than the feedforward entropy $\mathcal{T}_{X_{1}\to X_{1}}$. The mutual information, on the other hand, continues to obey its DPI, $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$. We expect further analytical work in the OU process to be able to support this empirical result.
@@
Acknowledgement We thank Manuel Reinhardt and Age Tjalma for useful discussions. This work is part of the Dutch Research Council (NWO) and was performed at the research institute AMOLF. This project has received funding from the European Research Council under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 885065).
Data availability A parallelized implementation of TE-PWS in Python is openly available on Zenodo at https://zenodo.org/doi/10.5281/zenodo.13617365.27.
Appendix A Appendix: Model details
@@
Figs. 2 a and b show transfer entropies in a minimal stochastic model for gene expression20, 28, 19. The model is a three-dimensional OU process with spring constants $a_{11}=\mu_{E}+\mu_{0}T_{\mu E}(T_{EG}-1)$, $a_{12}=-\mu_{0}(T_{EG}-1)$, $a_{13}=\mu_{0}T_{EG}$, $a_{21}=-T_{\mu E}[\beta_{G}-\mu_{E}-\mu_{0}T_{\mu E}(T_{EG}-1)]$, $a_{22}=-[\mu_{0}T_{\mu E}(T_{EG}-1)-\beta_{G}]$, $a_{23}=-[\beta_{G}-\beta_{\mu}-\mu_{0}T_{\mu E}T_{EG}]$, $a_{31}=a_{32}=0$ and $a_{33}=\beta_{\mu}$, and diffusion constants $D_{11}=D_{E}\mu_{0}^{2}$, $D_{22}=\beta_{\mu}\eta_{\mu}^{2}+\beta_{G}\eta_{G}^{2}+D_{E}\mu_{0}^{2}T_{\mu E}^{2}$, $D_{33}=\beta_{\mu}\eta_{\mu}^{2}$, $D_{12}=D_{21}=D_{E}\mu_{0}^{2}T_{\mu E}$, $D_{13}=D_{31}=0$ and $D_{23}=D_{32}=\beta_{\mu}\eta_{\mu}^{2}$, where the experimentally determined values of the parameters are $\mu_{0}=0.23h^{-1}$, $\beta_{\mu}=\beta_{G}=0.33h^{-1}$, $\beta_{E}=5.63h^{-}1$, $\eta_{E}=1.03$, $\eta_{\mu}=0.16$, $\eta_{G}=0.22$, $T_{EE}=1$, $T_{\mu E}=0.7$, $T_{EG}=1.3$, $\mu_{E}=\mu_{0}(1+T_{\mu E}-T_{EE})$ and $D_{E}=\eta_{E}^{2}/\beta_{E}$.20, 19.
@@
For Figs. 2c and d, the model is a two-dimensional OU process with parameters $a_{11}=1$, $a_{22}=2$, $a_{33}=1.5$, $a_{12}=a_{21}=0.5$, $a_{31}=0.2$, $a_{32}=0.3$, $D_{11}=1.5$, $D_{22}=0.5$, $D_{33}=1$ and $a_{13}=a_{23}=D_{12}=D_{23}=D_{13}=0$.
For models A and B in Fig. 3, we have chosen $a_{11}=a_{22}=1$, $a_{33}=-a_{32}=2$, $a_{13}=a_{31}=a_{23}=0$ and $D_{ij}=\delta_{ij}$.
@@
Error bars for all results are computed as twice the standard deviations in sets of statistically independent simulations.
Appendix B TE-PWS algorithm
@@
Central to TE-PWS is the computation of trajectory averages in a Monte-Carlo fashion over simulated trajectories, and the availability of trajectory probabilities on-the-fly in the full $d$-dimensional space. Thus for implementation in any stochastic model, two quantities need to be specified- a simulation method and the explicit functional form of the trajectory probability.
@@
For diffusion processes modeled by Eq. 1, we simulate it using an Euler-Maruyama scheme with a fixed small timestep. In our examples we have taken the timestep to be equal to the duration of trajectory segments for implementing the RR scheme, $\delta t$, for convenience. The propagation equation for the $(k+1)$-th step is $\mathbf{X}(k+1)=\mathbf{X}(k)+\delta t\mathbf{F}(k)+\sqrt{\delta t}\bm{\psi}(k)$, where $\bm{\psi}$ is a Gaussian random vector with mean zero and variance $\langle\psi_{i}(k)\psi_{j}(k^{{}^{\prime}})\rangle=2D_{ij}\delta_{k,k^{{}^{\prime}}}$. Making time discrete results in an $\mathcal{O}(\delta t)$ error which can be systematically made arbitrarily small. The probability density of the change of state $\Delta\mathbf{X}(k)$ can be written analytically through Ito discretization of the Onsager-Machlup action as 14, 15
@@
$\cdot\mathbf{D}^{-1}(\Delta\mathbf{X}(k)-\mathbf{F}(k)\delta t)/4\delta t\big{]}$ (12)
where $|\mathbf{D}|$ is the determinant of the diffusion constant matrix. This form also holds for systems with inertia if the generalized coordinate vector $\mathbf{X}$ contains both positions and velocities29, 30.
For jump processes, we simulate them with a Gillespie algorithm which is exact, i.e., does not make a timestep error. The Gillespie algorithm is implemented as follows. At any given time, the waiting time till the next jump is exponentially distributed. Hence at the start of the trajectory and after every jump, denoting the time as $t$, we sample a uniform random number $u\in[0,1]$ and determine the wait time $\tau$ by solving
@@
for $\tau$. Here $\lambda$ is the escape propensity from state $\mathbf{X}(t^{{}^{\prime}})$, given by the sum of all jump propensities taking the system out of the state $\mathbf{X}(t^{{}^{\prime}})$. When propagating all coordinates at once, $\lambda(t^{{}^{\prime}})$ is independent of $t^{{}^{\prime}}$ in between jumps, hence the equation is trivially solved for $\tau$. Then a second uniform random number is drawn to choose which jump to fire, with probabilities proportional to their jump propensities. The probability density of a trajectory segment $\mathbf{X}_{[t,t+\delta t]}$ is written as
@@
where $\alpha$ sums over all jumps, $N_{tot}$ in number.
@@
The TE-PWS algorithm for both diffusive and jump processes then proceeds as follows.
@@
  • Propagate $M_{1}$ trajectories of $X_{i}$ and $X_{j}$ in the full $d$-dimensional space (for diffusive processes, see Eq. 5). These trajectories are henceforth labeled with $(\nu)$.
  • For each pair of $(X_{i}^{(\nu)},X_{j}^{(\nu)})$ trajectories, propagate $M_{2}$ trajectories of hidden variables $X_{l}$ using the chosen reference distribution $P_{0}(X_{l,[0,k]})$ (for diffusive processes, the calculation here onwards corresponds to the computation of the numerator in Eq. 5).
@@
  • After every $\delta t$ time, recalculate logarithmic weights $w^{(\mu)}$ for the trajectories defined as the logarithm of the ratio between the joint distribution $P(X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{l,[0,k]})$. This ratio is proportional to the ratio between the conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ and the reference distribution $P_{0}(X_{l,[0,k]})$, as shown in the next section. Then calculate the uniformity in the weights with a uniformity parameter $\kappa=(\sum_{\mu}\exp w^{(\mu)})^{2}/\sum_{\mu}\exp(2w^{(\mu)})$, where sums of exponentials of weights are always performed with the Log-Sum-Exp trick31.
@@
  • Calculate the contribution to the transfer entropy in the space of $(X_{i},X_{j})$ at the $(k+1)$-th step, denoted as $\mathcal{T}_{a}^{(\nu)}[k]$. This is done by computing the average in Eq. 8 for Langevin processes, and taking an expectation of Eq. 9 for jump processes.
@@
  • If $\kappa<M_{2}/2$, resample the $M_{2}$ trajectories with the accumulated weights $w^{(\mu)}$. This means we sample $M_{2}$ trajectories randomly with weights $w^{(\mu)}$ from the simulated ensemble of $M_{2}$ trajectories, with replacement. Set all weights to zero after every resampling.
@@
  • Go back to the second item on this list and repeat the procedure for each $(X_{j}^{(\nu)})$ trajectory (for diffusive processes, this calculation corresponds to the computation of the denominator in Eq. 5). Specifically, propagate $M_{2}$ trajectories of hidden variables $(X_{i},X_{l})$ using the chosen reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$, accumulate weights for each trajectory as the ratio between the joint distribution $P(X_{i,[0,k]},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$, compute the contribution to the transfer entropy in the $(X_{j})$ space at the $(k+1)$-th step, denoted as $\mathcal{T}_{b}^{(\nu)}[k]$ (Eqs. 7 and 10), and resample if the uniformity in the weights is low, i.e., $\kappa<M_{2}/2$.
@@
  • Finally, compute $\mathcal{T}_{X_{i}\to X_{j}}$ using Eqs. 5 and 11 for Langevin and jump processes respectively.

### v2

@@ Abstract
We plot in Figs. 3c and d transfer entropy rates for the two models. Each data point is obtained from steady state trajectories using TE-PWS. In model A, increasing the feedback strength $f^{*}$ keeps $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ unchanged while $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$ monotonically increases. The former follows analytically from the expression for the transfer entropy in a two-dimensional OU process24 (see EM). The latter is an empirical result that arises because, when $f^{*}$ is increased, $X_{1}$ becomes increasingly correlated with $X_{2}$ while $X_{2}$ continues to be copied accurately by $X_{3}$. In contrast, in model B, both $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$ peak near the switching regime, phenomenologically similar to how a bistable three-node motif with feedback has been shown to behave through approximate theory15. The values of the transfer entropy rates in model B are also amplified many-fold compared to model A due to stronger correlations between $X_{1}$ and $X_{2}$ resulting from the stronger, nonlinear coupling, with the concomitant switching.
@@
In Fig. 3e we have plotted the amount of DPI violation in both models, quantified as the ratio of transfer entropy rates $\mathcal{T}^{*}\equiv\dot{\mathcal{T}}_{X_{1}\to X_{3}}/\dot{\mathcal{T}}_{X_{1}\to X_{2}}$, as a function of the ratio of the feedback to feedforward strengths, $f^{*}$. Surprisingly, we find that regardless of the nature of the variation of the individual transfer entropy rates with increasing feedback, i.e., monotonic or non-monotonic (Figs. 3c and d), the ratio $\mathcal{T}^{*}$ monotonically increases with increasing feedback (Fig. 3 e). Moreover, the ratio overcomes the DPI bound when the strength of the feedback becomes comparable to that of the feedforward coupling. Our results thus show that when the feedback $X_{2}\to X_{1}$ dominates over the feedforward interaction, $X_{1}\to X_{2}$, the feedforward entropy $\mathcal{T}_{X_{1}\to X_{3}}$ becomes larger than the feedforward entropy $\mathcal{T}_{X_{1}\to X_{2}}$. The mutual information, on the other hand, continues to obey its DPI, $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$. We expect further analytical work in the OU process to be able to support this empirical result.
@@
Acknowledgement We thank Manuel Reinhardt and Age Tjalma for useful discussions and Vahe Galstyan for a careful reading of the manuscript. This work is part of the Dutch Research Council (NWO) and was performed at the research institute AMOLF. This project has received funding from the European Research Council under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 885065).
Data availability Python implementations of TE-PWS for diffusive and jump processes, and data that reproduce the findings of this study, are openly available on Zenodo at https://zenodo.org/doi/10.5281/zenodo.1361736532.
Appendix A End Matter
@@
Model details. Figs. 2 a and b show transfer entropies in a minimal stochastic model for gene expression25, 11, 24. The model is a three-dimensional OU process with spring constants $a_{11}=\mu_{E}+\mu_{0}T_{\mu E}(T_{EG}-1)$, $a_{12}=-\mu_{0}(T_{EG}-1)$, $a_{13}=\mu_{0}T_{EG}$, $a_{21}=-T_{\mu E}[\beta_{G}-\mu_{E}-\mu_{0}T_{\mu E}(T_{EG}-1)]$, $a_{22}=-[\mu_{0}T_{\mu E}(T_{EG}-1)-\beta_{G}]$, $a_{23}=-[\beta_{G}-\beta_{\mu}-\mu_{0}T_{\mu E}T_{EG}]$, $a_{31}=a_{32}=0$ and $a_{33}=\beta_{\mu}$, and diffusion constants $D_{11}=D_{E}\mu_{0}^{2}$, $D_{22}=\beta_{\mu}\eta_{\mu}^{2}+\beta_{G}\eta_{G}^{2}+D_{E}\mu_{0}^{2}T_{\mu E}^{2}$, $D_{33}=\beta_{\mu}\eta_{\mu}^{2}$, $D_{12}=D_{21}=D_{E}\mu_{0}^{2}T_{\mu E}$, $D_{13}=D_{31}=0$ and $D_{23}=D_{32}=\beta_{\mu}\eta_{\mu}^{2}$, where the experimentally determined values of the parameters are $\mu_{0}=0.23h^{-1}$, $\beta_{\mu}=\beta_{G}=0.33h^{-1}$, $\beta_{E}=5.63h^{-}1$, $\eta_{E}=1.03$, $\eta_{\mu}=0.16$, $\eta_{G}=0.22$, $T_{EE}=1$, $T_{\mu E}=0.7$, $T_{EG}=1.3$, $\mu_{E}=\mu_{0}(1+T_{\mu E}-T_{EE})$ and $D_{E}=\eta_{E}^{2}/\beta_{E}$.25, 24. We also used $\delta t=9.2\times 10^{-5}\mu_{0}^{-1}$, $M_{1}=96$ and $M_{2}=200$.
@@
For Figs. 2c and d, the model is a two-dimensional OU process with parameters $a_{11}=1$, $a_{22}=2$, $a_{12}=a_{21}=0.5$, $D_{11}=1.5$, $D_{22}=0.5$ and $D_{12}=0$. Additionally, $\delta t=0.01a_{11}^{-1}$, $M_{1}=100$ and $M_{2}=1000$.
For models A and B in Fig. 3, we have chosen $a_{11}=a_{22}=1$, $a_{33}=-a_{32}=2$, $a_{13}=a_{31}=a_{23}=0$ and $D_{ij}=\delta_{ij}$. Additionally, $\delta t=0.01a_{11}^{-1}$, $M_{1}=100$ and $M_{2}=1000$.
@@
In all simulations in the main text, trajectories have been initialized from a histogram of the corresponding steady-state distributions. Error bars for all results are computed as twice the standard deviations using multiple sets of statistically independent simulations.
Analytical expression in two-dimensional OU process. In a two-dimensional OU process $(X_{1},X_{2})$ with spring constants $a_{ij}$ and diffusion constant constants $D_{ij}$ for $i,j\in\{1,2\}$, the steady-state transfer entropy rate $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ is given by24
$\dot{\mathcal{T}}_{X_{1}\to X_{2}}=\frac{1}{2}\left(r_{2}-a_{11}+\frac{D_{12}}{D_{22}}a_{21}\right)$ (12)
where $r_{2}=[a_{11}^{2}+(D_{11}/D_{22})a_{21}^{2}-2(D_{12}/D_{22})a_{11}a_{21}]^{1/2}$. For model A in the section Data processing inequality of the main text, the dynamics of $X_{1}$ and $X_{2}$ comprises that of a two-dimensional OU process because there is no feedback from $X_{3}$. Hence the transfer entropy rate $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ is given by Eq. 12, which demonstrably does not depend on the $X_{2}\to X_{1}$ feedback strength $a_{12}$, i.e., on $f^{*}$ (see Fig. 3c).
@@
TE-PWS algorithm. Central to TE-PWS is the computation of trajectory averages in a Monte-Carlo fashion over simulated trajectories, and the availability of trajectory probabilities on-the-fly in the full $d$-dimensional space. To implement TE-PWS for a given stochastic model, we need to specify the method to simulate the model and an explicit functional form of the trajectory probability.
@@
The diffusion processes modeled by Eq. 1 can be simulated using an Euler-Maruyama scheme with a fixed small timestep. In our examples we have taken the timestep to be equal to the duration of the trajectory segments for implementing the RR scheme, $\delta t$, for convenience. The propagation equation for the $(k+1)$-th step is $\mathbf{X}(k+1)=\mathbf{X}(k)+\delta t\mathbf{F}(k)+\sqrt{\delta t}\bm{\psi}(k)$, where $\bm{\psi}$ is a Gaussian random vector with zero mean and variance $\langle\psi_{i}(k)\psi_{j}(k^{{}^{\prime}})\rangle=2D_{ij}\delta_{k,k^{{}^{\prime}}}$. Making time discrete results in an $\mathcal{O}(\delta t)$ error which can be made arbitrarily small by systematically decreasing the timestep. The probability density of the change of state $\Delta\mathbf{X}(k)$ can be written analytically through Ito discretization of the Onsager-Machlup action as 20, 21
@@
$\qquad\cdot\mathbf{D}^{-1}(\Delta\mathbf{X}(k)-\mathbf{F}(k)\delta t)/4\delta t\big{]}$ (13)
where $|\mathbf{D}|$ is the determinant of the diffusion constant matrix. This form also holds for systems with inertia if the generalized coordinate vector $\mathbf{X}$ contains both positions and velocities33, 34.
Jump processes can be simulated with a Gillespie algorithm, which is exact, i.e., does not make a timestep error27. To clarify the quantities in the main text (Eqs. 9-11), and elucidate how they can be obtained in the Gillespie simulations, we briefly describe the scheme. At any given time, the waiting time till the next jump is exponentially distributed. Hence at the start of the trajectory and after every jump, denoting the time as $t$, we sample a uniform random number $u\in[0,1]$ and determine the wait time $\tau$ by solving
@@
for $\tau$. Here $\lambda(t^{{}^{\prime}})$ is the escape propensity from state $\mathbf{X}(t^{{}^{\prime}})$, given by the sum of the jump propensities $\mathcal{Q}_{\beta}(t^{{}^{\prime}})$ taking the system out of the state $\mathbf{X}(t^{{}^{\prime}})$ at time $t^{{}^{\prime}}$. When propagating all coordinates at once, $\lambda(t^{{}^{\prime}})$ is independent of $t^{{}^{\prime}}$ in between jumps; hence Eq. 14 is trivially solved for $\tau$. Then a second uniform random number is drawn to choose which jump to fire, with probabilities proportional to their jump propensities. The probability density of a trajectory segment $\mathbf{X}_{[t,t+\delta t]}$ is written as
@@
where $\alpha$ sums over all jumps in the full trajectory, $N_{\mathrm{tot}}$ in number, occurring at times $t_{\alpha}$.
@@
The elementary steps of the TE-PWS algorithm are the same for both diffusive and jump processes. In brief:
  1. 1.
@@
    Propagate $M_{1}$ trajectories of $X_{i}$ and $X_{j}$ in the full $d$-dimensional space. These trajectories are henceforth labeled with $(\nu)$.
    • Steps 2-5 are for computing the numerator in Eq. 5 for diffusive processes and the first term in Eq. 11 for jump processes.
  2. 2.
    For each pair of $(X_{i}^{(\nu)},X_{j}^{(\nu)})$ trajectories, propagate $M_{2}$ trajectories of hidden variables $X_{l}$ using the chosen reference distribution $P_{0}(X_{l,[0,k]})$. Initial conditions should be sampled from the same joint distribution $P(X_{i}^{(\nu)}(0),X_{j}^{(\nu)}(0),X_{l}(0))$ that $X_{i}^{(\nu)}(0)$ and $X_{j}^{(\nu)}(0)$ were drawn from.
  3. 3.
@@
    After every $\delta t$ time, recalculate logarithmic weights $w^{(\mu)}$ for the trajectories defined as the logarithm of the ratio between the joint distribution $P(X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{l,[0,k]})$, using Eqs. A and 14. This ratio is proportional to the ratio between the conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ and the reference distribution $P_{0}(X_{l,[0,k]})$, as proven in the SM. Then calculate the uniformity in the weights with a uniformity parameter $\kappa=(\sum_{\mu}\exp w^{(\mu)})^{2}/\sum_{\mu}\exp(2w^{(\mu)})$, where sums of exponentials of weights are always performed with the Log-Sum-Exp trick35.
  4. 4.
@@
    Calculate the contribution to the transfer entropy in the space of $(X_{i},X_{j})$ at the $(k+1)$-th step, denoted as $\mathcal{T}_{a}^{(\nu)}[k]$. This is done by computing the average in Eq. 8 for diffusive processes, and taking an expectation of Eq. 9 for jump processes (see SM).
  5. 5.
@@
    If $\kappa<M_{2}/2$, resample the $M_{2}$ trajectories with the accumulated weights $w^{(\mu)}$ (see Fig. 1c). This means we sample $M_{2}$ trajectories randomly with weights $w^{(\mu)}$ from the simulated trajectory ensemble, with replacement. Set all weights to zero after every resampling.
    • Step 6 is for computing the denominator in Eq. 5 for diffusive processes and the second term in Eq. 11 for jump processes.
  6. 6.
@@
    Akin to steps 2-5, for each $(X_{j}^{(\nu)})$ trajectory, propagate $M_{2}$ trajectories of $(X_{i},X_{l})$ using reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$; accumulate weights for each trajectory as the ratio between the joint distribution $P(X_{i,[0,k]},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$; compute the contribution to the transfer entropy in the $(X_{j})$ space at the $(k+1)$-th step, denoted as $\mathcal{T}_{b}^{(\nu)}[k]$ (Eqs. 7 and 10); and resample if the uniformity in the weights is low, i.e., $\kappa<M_{2}/2$.
  7. 7.
@@
    Finally, compute $\mathcal{T}_{X_{i}\to X_{j}}$ by combining $\mathcal{T}_{a}^{(\nu)}$ and $\mathcal{T}_{b}^{(\nu)}$ from all timesteps using Eqs. 5 and 11 for diffusive and jump processes respectively.
For clarity, a pseudocode to compute transfer entropy $\mathcal{T}_{X_{1}\to X_{3}}$ in a three-variable process is given in the SM. We note that TE-PWS is easily parallelized as it is a Monte-Carlo algorithm. Specifically, steps 1-7 can be executed independently for each of the $M_{1}$ trajectories mentioned in step 1, using parallel processors. The transfer entropy is computed at the end as an average over the results from all processors.
Supplemental Material:
Exact computation of Transfer Entropy with Path Weight Sampling
October 17, 2024
Appendix B Pseudocode